vidyo

New vidyo room:

Action Items:

News:

  • Lisa and Alessio phd thesis defense (2:30PM)

Framework & production

  • No news. Marco ran the demo in Commissione Calcolo e Reti in Rome last week. We'll ahve to hear from Marco to report on this
  • All users should check the dataset in LNL and give feedback on the ones that could be deleted [Federica will send around the gDoc]

Analysis

B physics

  • P'5 comments sent back to PLB

Charged Higgs to top-beauty

  • HIG conveners gave green light for unblinding, but Higgs Combination group asked to report on the cards and nuisances (Tue 27th).
  • Will unblind only the Pulls, and report to Higgs Combination before proceeding to a full unblinding

Dibosons, VZ resonance (V=W,Z)

  • Paul implemented all comments and prepared the paper for submission. Just waiting lastest checks from PubCom and button to be pushed for submitting it to JHEP.

4B full hadronic

Heavy New Quarks --> Heavy New Physics

  • Both 0PU and 200PU Neutrino samples are available, produced by UPSG
  • Working on 0PU to produce plots of physics quantities, will extend to 200PU and compare
  • On the Quark side, were stuck for a bug in the MC generator. News from Generator people

VH SM

  • No real news, activity still to get started. gDoc produced by Siew Yan circulated to Paul, Marco, Jacopo.

Same sign dileptons

EPR

  • Marina accepted to be remote FOG co-coordinator (--> 4 EPR months + shifts)
  • TSG looks for people to look at online and taking shifts (EPRs)

Trigger(less)

  • Andrea managed to integrate all firmware to read from TDCs, down to sending stream of triggerless data to PCIe
  • Code from Jacopo reads the stream and upacks it on the fly, producing occupancy plots. Some investigation will be needed (e.g. some "holes" in the occupancy plots from actually dead channels or mistakes in mapping, etc).
  • Will work on integrating code from Andrea and Sandro to get bx assignment on the fly and timebox plot

Detector

DT

Luminosity & Bril

  • Latest comparirson with full 2017 pixel showing some large discrepancies. Investigating.

Timing Layer

Tracker

A.O.B.:

-- Main.Franco Simonetto 04-10-2017 -- FrancoSimonetto - 2018-03-16

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2018-03-26 - JacopoPazzini
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback